\begin{thebibliography}{MMMXX}

\bibitem[AB24]{AB24}
K. Ardakov and L. Berger, 
\emph{Bounded functions on the character variety}, 
M\"{u}nster J. Math. (2024), to appear.

\bibitem[Box86]{Box86}
J.~Boxall, 
\emph{$p$-adic interpolation of logarithmic derivatives associated to certain Lubin-Tate formal groups}, 
Ann. Inst. Fourier \textbf{36} (1986), no.~3, 1--27.

\bibitem[Col79]{Col79}
R.~F. Coleman, 
\emph{Division values in local fields}, 
Invent. Math. \textbf{53} (1979), no.~2, 91--116.

\bibitem[Col02]{Col02}
P. Colmez, 
\emph{Espaces de Banach de dimension finie}, 
J. Inst. Math. Jussieu \textbf{1} (2002), 331--439.

\bibitem[Fon94]{Fon94}
J.-M. Fontaine
\emph{Le corps des p\'{e}riodes {$p$}-adiques},
Ast\'{e}risque (1994), no.~223, p.~59--111, With an appendix by Pierre
Colmez, P\'{e}riodes $p$-adiques (Bures-sur-Yvette, 1988).

\bibitem[Kat77]{Kat77}
N.~M. Katz, 
\emph{Formal groups and {$p$}-adic interpolation},
Journ\'{e}es {A}rithm\'{e}tiques de {C}aen, ({U}niv. {C}aen, 1976), 
Ast\'{e}risque No. 41--42, Soc. Math. France, Paris, 1977, pp.~55--65.

\bibitem[Kat81]{Kat81}
N.~M. Katz, 
\emph{Divisibilities, congruences, and {C}artier duality}, 
J. Fac. Sci. Univ. Tokyo Sect. IA Math. \textbf{28} (1981), no.~3, 667--678 (1982).

\bibitem[LT65]{LT65}
J. Lubin and J. Tate, 
\emph{Formal complex multiplication in local fields},
Ann. of Math. (2) \textbf{81} (1965), 380--387.

\bibitem[Tat67]{Tat67}
J.~T. Tate, 
\emph{{$p$}-divisible groups}, 
Proc. {C}onf. {L}ocal {F}ields ({D}riebergen, 1966), Springer, Berlin, 1967, pp.~158--183.

\end{thebibliography}

\documentclass[12pt,english,a4paper]{smfart}

\usepackage[T1]{fontenc}
\usepackage{lmodern}
\usepackage{smfthm}
\usepackage[headings]{fullpage}
\usepackage{amssymb}

\newcommand{\Hom}{\operatorname{Hom}}

\newcommand{\wotimes}{\operatorname{\widehat{\otimes}}}
\newcommand{\Gal}{\operatorname{Gal}}
\newcommand{\Qp}{\mathbf{Q}_p}
\newcommand{\Qpbar}{\overline{\mathbf{Q}}_p}
\newcommand{\Cp}{\mathbf{C}_p}
\newcommand{\QQ}{\mathbf{Q}}
\newcommand{\Gm}{\mathbf{G}_\mathrm{m}}
\newcommand{\calK}{\mathcal{K}}
\newcommand{\tph}{T_p H}
\newcommand{\tphx}{T_p^\times\! H}
\newcommand{\et}{\widetilde{\mathbf{E}}}
\newcommand{\at}{\widetilde{\mathbf{A}}}
\newcommand{\efont}{\mathbf{E}}
\newcommand{\afont}{\mathbf{A}}
\newcommand{\hatU}{\widehat{U}}
\newcommand{\OO}{\mathcal{O}}
\newcommand{\MM}{\mathfrak{m}}
\newcommand{\val}{\operatorname{val}}
\newcommand{\vp}{\val_p}
\newcommand{\Diff}{\operatorname{Diff}}
\newcommand{\pscal}[1]{\langle #1 \rangle}
\newcommand{\bigO}{\mathrm{O}}
\newcommand{\dcroc}[1]{[\![ #1 ]\!]}
\renewcommand{\phi}{\varphi}

\title{Galois measures and the Katz map}

\author{Laurent Berger}
\address{UMPA, ENS de Lyon \\ 
UMR 5669 du CNRS}
\email{laurent.berger@ens-lyon.fr}
\urladdr{https://perso.ens-lyon.fr/laurent.berger/}

\begin{document}

\begin{abstract}
The purpose of this paper is to explain the proofs of the results announced by Nick Katz in 1977, namely a description of ``Galois measures for Tate modules of height two formal groups over the ring of integers of a finite unramified extension of $\Qp$''. 
\end{abstract}

\date{\today}

\maketitle

\tableofcontents

\setlength{\baselineskip}{18pt}

\section*{Introduction}
\label{intro}

The purpose of this paper is to explain the proofs of the results announced by Nick Katz in 1977, namely a theory of ``Galois measures for Tate modules of height two formal groups over the ring of integers of a finite unramified extension of $\Qp$''. As Katz writes in the introduction of \cite{Kat81}, these results were announced in \cite{Kat77} but ``the details of this general theory remain unpublished''. 

In the author's joint paper \cite{AB24} with Konstantin Ardakov, Katz' results were proved in the case of a Lubin--Tate formal group attached to a uniformizer of $\QQ_{p^2}$ (theorem 1.6.1 of \cite{AB24}). In the present paper, we  simplify and clarify that proof, and extend it to all height two formal groups. We now describe our results in more detail. 

\subsection*{Formal groups}
Let $K$ be a finite unramified extension of $\Qp$ with ring of integers $\OO_K$ and residue field $k$. Let $\Qpbar$ be an algebraic closure of $K$, let $G_K = \Gal(\Qpbar/K)$ and let $\Cp$ be the completion of $\Qpbar$. Let $G$ be a formal group of dimension $1$ and height $2$ over $\OO_K$. Let $A(G) = \OO_K \dcroc{X}$ denote the coordinate ring of $G$. 

Let $H$ be the Cartier dual of $G$ (denoted by $G^{\vee}$ in \cite{Kat77}), so that (cf \cite{Tat67}) $H$ is also a formal group of dimension $1$ and height $2$. The Tate module $\tph$ parameterizes all formal group homomorphisms $G \to \Gm$ over $\OO_{\Cp}$. They are given by power series $t(X) \in \OO_{\Cp} \wotimes A(G) = \OO_{\Cp} \dcroc{X}$ such that $t(0)=1$ and $t(X \oplus_G Y) = t(X) \cdot t(Y)$. 

\subsection*{The covariant bialgebra of $G$}
Let $U(G)$ be the covariant bialgebra of $G$ (we use $U(G)$ instead of Katz' algebra $\Diff(G)$ of translation-invariant differential operators on $G$. We have $U(G) \simeq \Diff(G)$, see lemma \ref{ugdiff}). The set $U(G)$ is the set of all $\OO_K$-linear maps $A(G) \to \OO_K$ that vanish on some power of the augmentation ideal. If $f,g \in U(G)$, their product is defined by $(f \cdot g)(a(X)) = (f \otimes g)a(X \oplus_G Y)$ if $a(X) \in A(G)$. 

Let $\hatU(G)$ denote the set of $\OO_K$-linear maps $A(G) \to \OO_K$ that are continuous for the $(p,X)$-adic topology, so that $\hatU(G)$ is the $p$-adic completion of $U(G)$. Let $\pscal{\cdot,\cdot} : \hatU(G) \times A(G) \to \OO_K$ denote the evaluation pairing.

\subsection*{The Katz map}
Let $C^0_{\Gal}(\tph, \OO_{\Cp})$ denote the $\OO_K$-module of Galois continuous functions, namely those functions $f : \tph \to \OO_{\Cp}$ that are continuous and $G_K$-equivariant: $f(\sigma(t)) = \sigma(f(t))$ for all $t \in \tph$ and $\sigma \in G_K$. 

The evaluation pairing extends to $\pscal{\cdot,\cdot} : \hatU(G) \times (\OO_{\Cp} \wotimes A(G)) \to \OO_{\Cp}$. If $u \in \hatU(G)$, then $t \mapsto \pscal{u, t(X)}$ is a Galois continuous function $\tph \to \OO_{\Cp}$. The Katz map is
\[ \calK : \hatU(G) \to C^0_{\Gal}(\tph, \OO_{\Cp}), \]
defined by $\calK(u) (t) = \pscal{u, t(X)}$ (this is the map $(\ast)$ on page 59 of \cite{Kat77}). 

\subsection*{Galois measures}
Let $S$ be a $p$-adically complete and separated flat $\OO_K$-algebra. Applying the functor $\Hom_{\OO_K}(\cdot,S)$ to $\calK : \hatU(G) \to C^0_{\Gal}(\tph, \OO_{\Cp})$ gives an $S$-linear map (the map $(\ast\ast)$ on page 59 of \cite{Kat77})
\[ \calK^\ast : \Hom_{\OO_K} ( C^0_{\Gal}(\tph, \OO_{\Cp}), S) \to S \wotimes A(G). \]
If $f(X) \in S \wotimes A(G)$, let $\psi f \in S[1/p] \wotimes A(G)$ be defined by (compare with \S III of \cite{Col79}) 
\[ (\psi f)([p]_G(X)) = \frac{1}{p^2} \cdot \sum_{\pi \in G[p]} f(X \oplus_G \pi). \] 
We say that $f(X) \in S \wotimes A(G)$ is $\psi$-integral if $\psi^n f \in S \wotimes A(G)$ for all $n \geq 0$.
The main result claimed by Katz (see page 60 of \cite{Kat77}) is the following.

\begin{enonce*}{Theorem A}
\label{theoA}
The map $\calK^\ast$ is injective, and its image is the set of $\psi$-integral power series $f(X) \in S \wotimes A(G)$. 

Moreover, a Galois measure $\mu$ is supported on $\tphx$ if and only if $\psi(\calK^\ast(\mu)) = 0$.
\end{enonce*}

\subsection*{Main results}
The main result of this paper is a proof of theorem B below. Choose an element $t = (t_0,t_1,\hdots) \in \tphx = \tph \setminus p \cdot \tph$ and let $\calK_t : \hatU(G) \to \OO_{\Cp}$ denote the map $u \mapsto \calK(u)(t)$. Let $K_\infty$ denote the completion of the field generated over $K$ by the $t_n$ for $n \geq 1$. Since $\calK(u)$ is $G_K$-equivariant, the Ax-Sen-Tate theorem implies that $\calK(u)(t) \in \OO_{K_\infty}$ for all $t \in \tph$. 

\begin{enonce*}{Theorem B}
\label{theoB}
The map $\calK$ is injective, and the map $\calK_t : \hatU(G) \to \OO_{K_\infty}$ is surjective.
\end{enonce*}

Theorem A follows from theorem B by mostly formal arguments, that are carried out in detail in \cite{AB24} for the Lubin--Tate case, see in particular coro 3.4.10 of ibid (the hypothesis ``$\tau : G_L \to o_L^\times$ is surjective'' is replaced here by ``$G_K$ acts transitively on $\tphx$'', cf lemma \ref{galsurj} below). To keep this paper short, we focus on proving theorem B. 

Besides clarifying the arguments of \S 3 of \cite{AB24}, our proof also shows that $\hatU(G)$ has a perfectoid-like nature. In particular, the injectivity of $\calK$ is related to the following result in $p$-adic Hodge theory (cf 5.1.4 of \cite{Fon94}, and its Lubin--Tate generalizations such as prop 9.6 of \cite{Col02}): $\{ x \in \at^+$ such that $\theta \circ \phi^n(x)=0$ for all $n \geq 0\} = \pi \cdot \at^+$.

\section{Preliminaries}

In all this paper, we let $q=p^2$. The multiplication-by-$p$ map on $G$ is given by a power series $[p]_G(X)  = \sum_{i \geq 1} r_i X^i$ with $r_1=p$ and $r_i \in p \OO_K$ for $i \leq q-1$ and $r_q \in \OO_K^\times$.

\begin{lemm}
\label{galsurj}
If $t,u \in \tphx$, there exists $\sigma \in G_K$ such that $u = \sigma(t)$.
\end{lemm}

\begin{proof}
For all $n \geq 0$, in the Weierstrass factorization of $[p^{n+1}]_H(X) / [p^n]_H(X)$, the distinguished poly\-nomial is Eisenstein, as its constant term is $p$ and $K/\Qp$ is unramified. 

Its roots are therefore conjugate by $G_K$.
\end{proof}

\begin{lemm}
\label{chgvar}
There is a change of variables such that $[p]_G(X) = pX+ \bigO(X^q)$.
\end{lemm}

\begin{proof}
If $r(X) = \sum_{i =1}^{q-1} r_i X^i$, then $r(X)+X^q$ is a Lubin--Tate power series for the uniformizer $p$, so by lemma 1 of \cite{LT65} there exists a reversible power series $h(X)$ such that $h^{-1}  \circ (r(X)+X^q) \circ h = pX+X^q$. We then have $h^{-1}  \circ [p]_G(X) \circ h = pX+\bigO(X^q)$.
\end{proof}

A coordinate satisying the above conditions is said to be clean. Let $\log_G(X) \in K \dcroc{X}$ denote the logarithm attached to the formal group $G$.

\begin{lemm}
\label{logspec}
If $X$ is clean, then $\log_G(X) = X + \alpha \cdot X^q/p + \bigO(X^{q+1})$ for some $\alpha \in \OO_K^\times$. 
\end{lemm}

Let $t$ be an element of $\tph = \Hom_{\OO_{\Cp}}(G,\Gm)$, and let $t(X) = 1 + \sum_{i \geq 1} a_i(t) X^i \in \OO_{\Cp} \dcroc{X}$ be the corresponding Hom from $G$ to $\Gm$. 

\begin{lemm}
\label{galth}
If $\sigma \in G_K$ then $a_n(\sigma(t)) = \sigma(a_n(t))$ for all $n \geq 1$.
\end{lemm}

\begin{proof}
The isomorphism $\tph = \Hom_{\OO_{\Cp}}(G,\Gm)$ is $G_K$-equivariant.
\end{proof}

\begin{lemm}
\label{acoeff}
We have $a_n(t) = a_1(t)^n/n!$ if $n \leq q-1$ and $a_q(t) = a_1(t)^q/q! + \alpha \cdot a_1(t)/p$.
\end{lemm}

\begin{proof}
Since $t(X) = \exp( a_1(t) \cdot \log_G(X) )$, the claim follows from lemma \ref{logspec}.
\end{proof}

\begin{prop}
\label{katzval}
If $t \in \tphx$,  the abscissa of the breakpoints of the Newton polygon of $t(X)-1$ are the $p^m$ with $m \geq 0$, and $\vp(a_{p^m}(t)) = 1/p^{m-1}(q-1)$ for all $m \geq 0$.
\end{prop}

\begin{proof}
The Newton polygon of $t(X)-1$ is independant of the choice of coordinate $X$, and we choose a clean coordinate on $G$.
We first prove that $\vp(a_1(t)) \leq p/(q-1)$. Let $\zeta$ be a primitive $p$-th root of $1$. Since $t \in \tphx$, there exists $\eta \in G[p]$ such that $t(\eta)=\zeta$. We have $\vp(\eta)=1/(q-1)$ and $\vp(\zeta-1)=1/(p-1)$. 
Since $X$ is a clean coordinate on $G$, we have $a_n(t) = a_1(t)^n/n!$ if $n \leq q-1$ by lemma \ref{acoeff}.
If $\vp(a_1) > p/(q-1)$, then for $n < q$
\[ \vp(a_n(t) \eta^n) = \vp\left(a_1(t)^n \frac{\eta^n}{n!}\right)  > \frac{np+n}{q-1}-\frac{n-s_p(n)}{p-1} = \frac{s_p(n)}{p-1}  \geq \frac{1}{p-1}, \] 
while $\vp(\eta^n) = n/(q-1) > 1$ if $n \geq q$. It is therefore not possible to have $t(\eta)-1=\zeta-1$, so that $\vp(a_1(t)) \leq p/(q-1)$.

For every $m \geq 1$, $t(X)-1$ has $p^m-p^{m-1}$ zeroes in $G[p^m] \setminus G[p^{m-1}]$ and they are all of valuation $1/(q^m-q^{m-1})$. Since $\sum_{m \geq 1} (p^m-p^{m-1}) / (q^m-q^{m-1}) = p/(q-1)$, the theory of Newton polygons tells us that $\vp(a_1(t)) = p/(q-1)$, and then that $t(X)-1$ cannot have other zeroes. 

The valuations of the $a_{p^m}(t)$ can then be read off the Newton polygon of $t(X)-1$.
\end{proof}

\begin{rema}
\label{kalm2}
{\hspace{1pt}}
\begin{enumerate}
\item Compare with coro 3.5.8 of \cite{AB24};
\item Compare with the proof of lemma 13 of \cite{Box86};
\item Prop \ref{katzval} applied to $m=0$ gives us lemma 1 on page 62 of \cite{Kat77};
\item Lemma \ref{acoeff} and the fact that $\vp(a_q(t)) = 1/p(q-1)$ imply that $\vp(a_1(t) \cdot ( a_1(t)^{q-1}/q! + \alpha /p )) = 1/p(q-1)$
and hence $\vp(a_1(t)^{q-1} p / \alpha q! +1) = 1 - 1/p$, compare with lemma 2 on page 63 of \cite{Kat77}.
\end{enumerate}
\end{rema}

\section{The Katz map}

Let $C^0_{\Gal}(\tph, \OO_{\Cp})$ denote the set of Galois continuous functions $f : \tph \to \OO_{\Cp}$, namely those maps that are continuous and $G_K$-equivariant: $f(\sigma(t)) = \sigma(f(t))$ for all $t \in \tph$ and $\sigma \in G_K$. By lemma \ref{galsurj}, any two elements of $p^n \cdot \tphx$ are Galois conjugates, so if we fix $t \in \tphx$, a function $f \in C^0_{\Gal}(\tph, \OO_{\Cp})$ is determined by $\{ f(p^n t) \}_{n \geq 0}$. In addition, we have $f(p^n t) \to f(0)$ as $n \to +\infty$ by continuity, and $f(0) \in \OO_K$ by the Ax-Sen-Tate theorem since $f(0)$ is fixed by $G_K$. 

Write $t = (t_0,t_1, \hdots) \in \tph$ and let $K_n$ denote the extension of $K$ generated by $t_n$ and let $K_\infty$ be the completion of $\cup_{n \geq 0} K_n$. By the Ax-Sen-Tate theorem, we have $f(p^n t) \in \OO_{K_\infty}$ for all $n$. Let $\prod_{n \geq 0}' \OO_{K_\infty}$ denote the $\OO_K$-algebra of sequences $\{ f_n \}_{n \geq 0}$ with $f_n\in \OO_{K_\infty}$ and such that $\{ f_n \}_{n \geq 0}$ converges to an element of $\OO_K$ as $n \to +\infty$. 
The map $f \mapsto \{ f(p^n t) \}_{n \geq 0}$ gives an isomorphism $C^0_{\Gal}(\tph, \OO_{\Cp}) \to \prod_{n \geq 0}' \OO_{K_\infty}$. 

Recall that $U(G)$ is the covariant bialgebra of $G$, namely the set of all $\OO_K$-linear maps $A(G) \to \OO_K$ that vanish on some power of the augmentation ideal. If $f,g \in U(G)$, their product is defined by $(f \cdot g)(a(X)) = (f \otimes g)a(X \oplus_G Y)$.  In \cite{Kat77} and \cite{Kat81}, Katz introduces the algebra $\Diff(G)$ of translation-invariant differential operators on $G$.

\begin{lemm}
\label{ugdiff}
If $D \in \Diff(G)$, then $[ f \mapsto D(f)(0) ] \in U(G)$, and this map gives rise to an isomorphism between the $\OO_K$-algebras $\Diff(G)$ and $U(G)$.
\end{lemm}

\begin{proof}
The proof is the same as that in \S 3.2 of \cite{AB24} for the Lubin--Tate case: for $n \geq 0$, let $u_n \in U(G)$ be the map $u_n : A(G) \to \OO_K$ given by $u_n(\sum_{i \geq 0} b_i X^i)=b_n$. We have $u_n \cdot u_m = \sum_{k \geq 0} s_{k,n,m} u_k$ where $(X \oplus_G Y)^k = \sum_{n,m \geq 0} s_{k,n,m} X^n Y^m$ for $k \geq 0$, and we get the same structure constants as in $\Diff(G)$, see for instance (1.2) of \cite{Kat81}.
\end{proof}

Recall that $\hatU(G)$ denotes the set of $\OO_K$-linear maps $A(G) \to \OO_K$ that are continuous for the $(p,X)$-adic topology, so that $\hatU(G)$ is the $p$-adic completion of $U(G)$. The evaluation pairing $\pscal{\cdot,\cdot} : \hatU(G) \times A(G) \to \OO_K$  extends to $\pscal{\cdot,\cdot} : \hatU(G) \times (\OO_{\Cp} \wotimes A(G)) \to \OO_{\Cp}$.

\begin{defi}
\label{defkatz}
The Katz map is the map $\calK : \hatU(G) \to C^0_{\Gal}(\tph, \OO_{\Cp})$ defined by $\calK(u) (t) = \pscal{u, t(X)}$. 
\end{defi}

The map $\calK(u) : \tph \to \OO_{\Cp}$ is Galois continuous by lemma \ref{galth}. The Katz map $\calK$ is an $\OO_K$-algebra homomorphism (the proof is the same as that of lemma 3.3.5 of \cite{AB24}). Let $\phi_C : C^0_{\Gal}(\tph, \OO_{\Cp}) \to C^0_{\Gal}(\tph, \OO_{\Cp})$ denote the map given by $(\phi_C f)(t) = f(pt)$. We have a map $U(\phi_G) : \hatU(G) \to \hatU(G)$ coming by duality from the map $\phi_G : A(G) \to A(G)$ given by $a(X) \mapsto a([p]_G(X))$.

\begin{lemm}
\label{phikatz}
We have $\calK \circ U(\phi_G) = \phi_C \circ \calK$.
\end{lemm}

\begin{proof}
The proof is the same as that of lemma 3.3.6 of \cite{AB24}, where $U(\phi_G)$ is denoted by $\phi^*$.
\end{proof}

\begin{rema}
\label{bprime}
Since $\hatU(G) = \{ \sum_{n \geq 0} \lambda_n u_n$ with $\{ \lambda_n\}_{n \geq 0} \in c^0(\OO_K) \}$, and $u_n(t(X))= a_n(t)$,
we can reformulate theorem B as follows. 
\begin{enumerate}
\item If $t \in \tphx$,  every $x \in \OO_{K_\infty}$ can be written as $x = \sum_{n \geq 0} \lambda_n a_n(t)$ with $\{ \lambda_n\} \in c^0(\OO_K)$;
\item If $\{ \lambda_n\} \in c^0(\OO_K)$ and $\sum_{n \geq 0} \lambda_n a_n(t) = 0$ for all $t \in \tph$, then $\lambda_n = 0$ for all $n$.
\end{enumerate}
\end{rema}

\section{Surjectivity}

We now prove that $\calK_t : \hatU(G) \to \OO_{K_\infty}$ is surjective if $t = (t_0,t_1, \hdots) \in \tphx$. 

For $n \geq 0$, let $G_n=G[p^n]$ and $H_n=H[p^n]$ and recall that $K_n = K(t_n)$. The inclusion $G_n \to G_{n+1}$ is (\S 2.3 of \cite{Tat67}) the Cartier dual of $[p]_H : H_{n+1} \to H_n$ (and vice versa). We have $A(G_n) = \OO_K\dcroc{X}/\phi_G^n(X)$. Let $U(G_n) = \Hom_{\OO_K}(A(G_n),\OO_K) \subset \hatU(G)$. Cartier duality gives an isomorphism $U(G_n) = A(H_n)$, so that $U(G_n) = \OO_K\dcroc{Y_n}/\phi_H^n(Y_n)$. On $U(G_n)$, we have $U([p]_G) = [p]_H$ so that $U(\phi_G) = \phi_H$. The natural inclusion $U(G_n) \to U(G_{n+1})$ is the map $A(H_n) \to A(H_{n+1})$ that comes from $[p]_H : H_{n+1} \to H_n$. Its image is $\phi_H(U(G_{n+1}))$. 
 If $U(G_n)_K = K \otimes_{\OO_K} U(G_n)$, then 
\[ U(G_n)_K =  K \otimes_{\OO_K} \OO_K\dcroc{Y_n}/\phi_H^n(Y_n) = K_n \times K_{n-1} \times \cdots \times K_1 \times K_0. \] 

Fix $t \in \tphx$. For $n \geq 0$, let $\kappa_n : \hatU(G) \to \OO_{K_\infty}$ denote the map $u \mapsto \calK(u)(p^n t)$. Lemma \ref{phikatz} implies that $\kappa_n(\phi_H(u)) = \kappa_{n+1}(u)$ if $u \in U(G_m)$ for some $m, n \geq 0$.

\begin{prop}
\label{katsurj}
For all $n \geq 1$, the image of the map $\kappa_0 : U(G_n)_K \to K_\infty$ is $K_n$.
\end{prop}

\begin{proof}
We first prove that the image of $\kappa_0 : U(G_n)_K \to K_\infty$ is contained in $K_n$. By Cartier duality, we have $H_n = \Hom_{\OO_{\Cp}}(G_n,\Gm)$, and the map $t_n(X) \in \Hom_{\OO_{\Cp}}(G_n,\Gm)$ corresponding to $t_n$ is given by $t(X) \bmod{\phi_G^n(X)}$. We then have 
\[ t_n(X) \in \OO_{\Cp} \otimes A(G_n) = \Hom_{\OO_{\Cp}} ( \OO_{\Cp} \otimes U(G_n) , \OO_{\Cp}) \]
The resulting map $U(G_n)_K \to \Cp$ is given by $u \mapsto u(t(X) \bmod{\phi_G^n(X)}) = u(t)$, namely the restriction of $\kappa_0$ to $U(G_n)_K$. For all $\sigma \in G_{K_n}$, $\sigma(t_n) = t_n$ and hence $\sigma(t_n(X)) = t_n(X)$ so that $\sigma(\kappa_0(u)) = \kappa_0(u)$ if $u \in U(G_n)_K$. By the Ax-Sen-Tate theorem, $\kappa_0(u) \in K_n$.

We now prove that the image of $\kappa_0$ is $K_n$. Let $[p^n]_G(X) = f(X) u(X)$ be the Weierstrass factorization of $[p^n]_G(X)$, where $f(X)$ is a distinguished polynomial of degree $q^n$ and $u(X)$ is a unit. Thanks to Weierstrass division by $f(X)$, we can write 
\[ \OO_K \dcroc{X} = (\oplus_{i=0}^{q^n-1} \OO_K X^i) \oplus f(X) \OO_K \dcroc{X}. \]
For $0 \leq r \leq q^n-1$, let $w_r \in U(G_n)_K$ be the $\OO_K$-linear form that maps $X^i$ to $\delta_{ir}$ for $i \leq q^n-1$ and maps $f(X) \OO_K \dcroc{X}$ to $0$. It maps $X^{j+q^n}$ to $p \OO_K$ for all $j \geq 0$ since $f(X) = f_{q^n} X^{q^n} +p g(X)$ with $f_{q^n} \in \OO_K^\times$ and hence $f_{q^n} X^{j+q^n} = X^j f(X) - p X^j g(X)$. 

By prop \ref{katzval}, we have $\vp(a_{p^m}(t)) = 1/p^{m-1}(q-1)$ for all $m \geq 0$. In particular, setting $m=2n-1$, so that $p^m=q^n/p \leq q^n-1$, we have  $\vp(a_{p^m}(t)) = 1/q^{n-1}(q-1)$. We have \[ \kappa_0(w_{p^{2n-1}}) = w_{p^{2n-1}}(t(X)) \equiv a_{p^m}(t) \bmod{p}, \] so that $\kappa_0(U(G_n)_K)$ contains a uniformizer of $K_n$ and is therefore equal to $K_n$.
\end{proof}

\begin{rema}
\label{surjab}
Compare with (the proof of) prop 3.6.7 of \cite{AB24}.
\end{rema}

\begin{prop}
\label{katzum}
Take $n \geq 1$ and fix an isomorphism $U(G_n) = \OO_K\dcroc{Y}/\phi_H^n(Y)$. 

There exists $(u_n,u_{n-1},\hdots,u_1,u_0)$ with $u_k \in K_\infty$ and $[p]_H(u_k) = u_{k-1}$ for $1 \leq k \leq n$ and $u_1 \neq 0$ and $u_0=0$ such that the map $\kappa_j : U(G_n)_K \to K_\infty$ is given by $P(Y) \mapsto P(u_{n-j})$ if $j \leq n$, and by $P(Y) \mapsto P(0)$ if $j \geq n$.
\end{prop}

\begin{proof}
For all $j$, there exists a root $u_{n-j}$ of $\phi_H^n(Y)$ in $\MM_{K_\infty}$ such that the $K$-algebra map $\kappa_j : K \otimes_{\OO_K} \OO_K\dcroc{Y}/\phi_H^n(Y) \to K_\infty$ is given by $P(Y) \mapsto P(u_{n-j})$. In addition, we have $\kappa_{j-1} ( \phi_H(Y) ) = \kappa_j(Y)$ so that $[p]_H(u_{n-j+1}) = u_{n-j}$. 

We have $\kappa_n(Y) = \kappa_0 ( \phi_H^n(Y) ) = 0$ so that $u_0=0$, and finally since $\kappa_0 : U(G_1)_K \to K_1$ is surjective by prop \ref{katsurj}, and 
\[ \kappa_{n-1}(U(G_n)_K) = \phi_C^{n-1} \circ \kappa_0(U(G_n)_K) = \kappa_0(\phi_H^{n-1}(U(G_n))_K) = \kappa_0(U(G_1)_K), \]
the element $u_1$ generates $K_1$ over $K$ and so $u_1 \neq 0$.
\end{proof}

\begin{coro}
\label{katsurjint}
The map $\kappa_0 : U(G_n) \to \OO_{K_n}$ is surjective for all $n \geq 0$.
\end{coro}

\begin{proof}
The map $\kappa_0 : U(G_n) \to \OO_{K_\infty}$ is given by $P(Y_n) \mapsto P(u_n)$ as in prop \ref{katzum}, so that its image is $\OO_{K_n}$ since $u_n$ is a uniformizer of $K_n$.
\end{proof}

\begin{coro}
\label{kappasurj}
The map $\kappa_0 : \hatU(G) \to \OO_{K_\infty}$ is surjective.
\end{coro}

This proves the surjectivity part of Theorem B since $\kappa_0 = \calK_t$.

\section{Injectivity}

We now prove the injectivity of the Katz map $\calK : \hatU(G) \to C^0_{\Gal}(\tph,\OO_{\Cp})$. Recall that $U(G_n) = \Hom_{\OO_K}(\OO_K\dcroc{X}/\phi_G^n(X),\OO_K)$ and that we have an injection $U(G_n) \to \hatU(G)$.

\begin{lemm}
\label{undense}
The $\OO_K$-module $\cup_{n \geq 1} U(G_n)$ is $p$-adically dense in $\hatU(G)$.
\end{lemm}

\begin{proof}
Take $k,d \geq 0$ and $u \in \hatU(G)$ such that $u(X^d \OO_K \dcroc{X}) \subset p^k \OO_K$, and  $n \geq 1$ such that $[p^n]_G(X) \in (X^d,p^k)$. Let $[p^n]_G(X) = f(X) u(X)$ be the Weierstrass factorization of $[p^n]_G(X)$, where $f(X)$ is distinguished of degree $q^n$ and $u(X)$ is a unit. Write $\OO_K \dcroc{X} = (\oplus_{i=0}^{q^n-1} \OO_K X^i) \oplus f(X) \OO_K \dcroc{X}$. Define $w \in U(G_n)$ by $w=u$ on $1,X,\hdots,X^{q^n-1}$ and $w=0$ on $f(X) \OO_K \dcroc{X}$. If $j \geq 0$, then $X^{j+q^n} \in f(X) \OO_K \dcroc{X} +(\oplus_{i=d+j}^{q^n+j-1} \OO_K X^i)  + p^k \OO_K \dcroc{X}$. By induction on $j$, we have $w(X^{j+q^n}) \in p^k \OO_K$ for all $j \geq 0$ and therefore $w-u \in p^k \hatU(G)$.
\end{proof}

The group $G_K$ acts transitively on $\tphx$ by lemma \ref{galsurj}, so that in the isomorphisms $U(G_n) = \OO_K\dcroc{Y_n}/\phi_H^n(Y_n)$, it is possible to change the coordinates $Y_n$ to get a common $u =(\hdots,u_1,u_0)$ for all $U(G_n)_K$ in prop \ref{katzum}. We can then write $U(G_n) = \OO_K \dcroc{ \phi_H^{-n}(Y) } / Y$ for each $n$, with each transition map $U(G_{n-1}) \to U(G_n)$ sending $Y$ to $Y$ and the map $\kappa_j$ sending $\phi_H^{-n}(Y)$ to $u_{n-j}$. Lemma \ref{undense} and the fact that $U(G_n) \cap p \cdot \hatU(G) = p \cdot U(G_n)$ for all $n \geq 0$ imply that $\hatU(G)$ is the $p$-adic completion of $\cup_{n \geq 0} \OO_K \dcroc { \phi_H^{-n}(Y) } / Y$.

Let $\afont$ be the $p$-adic completion of $\cup_{n \geq 0} \OO_K \dcroc { \phi_H^{-n}(Y) }$ and let $\theta : \afont \to \OO_{K_\infty}$ be the $\OO_K$-linear ring homomorphism that sends $\phi_H^{-n}(Y)$ to $u_n$ so that $\hatU(G) = \afont/Y$ and $\kappa_0 = \theta$. 

Let $\efont$ denote the ring $\cup_{n \geq 0} k \dcroc{ \phi_H^{-n}(Y) }$. The valuation $\val_Y$ is compatibly defined on each $k \dcroc{ \phi_H^{-n}(Y) }$ and hence on $\efont$ (indeed, $\phi_H(Y) \in Y^q \cdot k \dcroc{Y}^\times$ so that if $x \in k \dcroc{Y_n} \subset  k \dcroc{Y_{n+1}}$ with $Y_n = \phi_H(Y_{n+1})$, then $\val_{Y_{n+1}}(x) = q \cdot \val_{Y_n}(x)$).

\begin{lemm}
\label{divy}
If $x \in \efont$, then $x \in Y \cdot \efont$ if and only if $\val_Y(x) \geq 1$.
\end{lemm}

Let $\theta : \efont \to \OO_{K_\infty} / p$ be the $k$-linear ring homomorphism that sends $\phi_H^{-n}(Y)$ to $u_n$ so that $\efont = \afont / p \afont$ with compatible $\theta$.

\begin{lemm}
\label{kerthetet}
We have $\ker(\theta : \efont \to \OO_{K_\infty} / p) = Y/\phi_H^{-1}(Y) \cdot \efont$. 
\end{lemm}

\begin{proof}
It is enough to prove that for all $n \geq 1$, $\ker(\theta : k \dcroc{ \phi_H^{-n}(Y) } \to \OO_{K_\infty} / p)$ is generated by $Y/\phi_H^{-1}(Y)$. Let $Q_n(X)$ be the minimal polynomial of $u_n$ over $K$, so that $Q_n(X) \in \OO_K[X]$ is monic of degree $q^{n-1}(q-1)$. If $P(X) \in \OO_K \dcroc{X}$, we can write it as $P=SQ_n+R$ with $\deg R < q^{n-1}(q-1)$. If $P(u_n) \in p \cdot \OO_{K_n}$, then $R(u_n) \in p \cdot \OO_{K_n}$ and since $\vp(u_n) = 1/q^{n-1}(q-1)$, this implies that $R(X) \in p \cdot \OO_K[X]$. 
The claim now follows from this, and from the fact that $\theta(\phi_H^{-n}(Y)) = u_n$ so that $\ker(\theta)$ is generated by $\phi_H^{-n}(Y)^{q^{n-1}(q-1)}$ and hence by $Y/\phi_H^{-1}(Y)$ since $\phi_H(Y) = Y^q \cdot f(Y)$ with $f(Y) \in k \dcroc{Y}^\times$.
\end{proof}

\begin{prop}
\label{kerthet}
We have $\ker(\theta : \afont \to \OO_{K_\infty}) = Y/\phi_H^{-1}(Y) \cdot \afont$. 
\end{prop}

\begin{proof}
If $x \in \ker(\theta)$, then $\overline{x} \in \efont$ is also killed by $\theta$ and is  divisible by $Y/\phi_H^{-1}(Y)$ in $\efont$ by lemma \ref{kerthetet}. We can therefore write $x= Y/\phi_H^{-1}(Y) \cdot x_1 + p y_1$ with $x_1,y_1 \in \afont$ and $\theta(y_1)=0$. By induction, we can write $x= Y/\phi_H^{-1}(Y) \cdot x_k + p^k y_k$ with $x_k,y_k \in \afont$ and $\theta(y_k)=0$ for all $k \geq 1$. Since $\afont$ is $p$-adically complete, this implies the claim.
\end{proof}

\begin{prop}
\label{kerkatz}
We have $\{ x \in \afont$, $\theta \circ \phi_H^n(x)=0$ for all $n \geq 0\} = Y \cdot \afont$.
\end{prop}

\begin{proof}
One inclusion is clear, since $\theta \circ \phi_H^n(Y)=0$ for all $n \geq 0$. We now prove the reverse inclusion. Prop \ref{kerthet} shows that $\ker(\theta) = Y/\phi_H^{-1}(Y) \cdot \afont$ and therefore that for all $j \geq 0$, $\ker(\theta \circ \phi_H^j) = \phi_H^{-j}(Y)/\phi_H^{-j-1}(Y) \cdot \afont$. For $n \geq 0$, let $I_n$ denote the set of $x \in \afont$ such that $(\theta \circ \phi_H^j) (x) = 0$ for $0 \leq j \leq n$. Since $(\theta \circ \phi_H^\ell)(\phi_H^{-j}(Y)/\phi_H^{-j-1}(Y)) \neq 0$ if $\ell < j$, and $Y/\phi_H^{-1}(Y) \cdot \phi_H^{-1}(Y)/\phi_H^{-2}(Y) \cdots \phi_H^{-(n-1)}(Y)/\phi_H^{-n}(Y) = Y/\phi_H^{-n}(Y)$, we have $I_n =  Y/\phi_H^{-n}(Y) \cdot \afont$. Let $I = \cap_{n \geq 0} I_n = \{ x \in \afont$, $\theta \circ \phi_H^n(x)=0$ for all $n \geq 0\}$.

The above reasoning and lemma \ref{divy} imply that in $\efont$, we have $\overline{I}= Y \cdot \efont$. Hence if $x \in I$, then $\overline{x} \in Y \cdot \efont$. We can therefore write $x= Y \cdot x_1 + p y_1$ with $x_1 \in \afont$ and $y_1 \in I$. By induction, we can write $x= Y \cdot x_k + p^k y_k$ with $x_k \in \afont$ and $y_k \in I$ for all $k \geq 1$. Since $\afont$ is $p$-adically complete, this implies the claim.
\end{proof}

\begin{rema}
\label{colthet}
Compare with prop 9.6 of \cite{Col02} (and 5.1.4 of \cite{Fon94}), noting however that our $\efont$ is not complete for the $Y$-adic topology. We could actually replace $\efont$ and $\afont$ above with their $Y$-adic completions, since in any case $\hatU(G) = \afont/Y$. In the Lubin--Tate case, we would then have $\efont = \et^+_K$ and $\afont = \at^+_K$ in the notation of ibid.

In general, $\phi_H(Y) = f(Y^q)$ in $k \dcroc{Y}$ for some reversible $f(Y)$ so that $\efont$ is still perfect.
\end{rema}

\begin{coro}
\label{katzinj}
The map $\calK : \hatU(G) \to C^0_{\Gal}(\tph,\OO_{\Cp})$ is injective.
\end{coro}

\begin{proof}
If we write $C^0_{\Gal}(\tph,\OO_{\Cp}) = \prod'_{n \geq 0} \OO_{K_\infty}$, then $\calK : \hatU(G) \to C^0_{\Gal}(\tph,\OO_{\Cp})$ comes from $\{ \theta \circ \phi_H^n \}_{n \geq 0} : \afont \to \prod'_{n \geq 0} \OO_{K_\infty}$ and the claim results from prop \ref{kerkatz} above and the fact that $\hatU(G) = \afont/Y$ by lemma \ref{undense}.
\end{proof}

This finishes the proof of Theorem B.

\end{document}